# A Variational Principle for Extended Irreversible Thermodynamics: Heat Conducting Viscous Fluids

### v1

@@ 2.2 Classical irreversible thermodynamics
Variational formulation in the material description.
Besides the Lagrangian, in the irreversible case one needs to introduce the thermodynamic fluxes associated with the irreversible processes considered. In the case of viscosity and heat conduction the corresponding fluxes in the material description are given by the viscous Piola-Kirchhoff tensor $\boldsymbol{P}$ and the material entropy flux $\boldsymbol{J}$. In a similar way to the finite dimensional case recalled in (1.7)–(1.9), the variational formulation needs the introduction of the thermal displacement $\Gamma(t,\mathbf{X})$ whose time rate of change is identified with the temperature in the material frame, i.e., $\frac{d}{dt}\Gamma(t,\mathbf{X})=\mathfrak{T}(t,\mathbf{X})$, as well as the additional entropy density variable $\Sigma(t,\mathbf{X})$ whose time rate of change is identified with the rate of internal entropy production of the continua $\frac{d}{dt}\Sigma(t,\mathbf{X})=\mathfrak{I}(t,\mathbf{X})\geq 0$.
@@
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\frac{d}{dt}\frac{\partial\mathfrak{L}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\mathfrak{L}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\mathfrak{L}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\mathfrak{L}}{\partial S}(\dot{S}+\operatorname{DIV}\boldsymbol{J})=\boldsymbol{P}:\nabla\dot{\boldsymbol{\varphi}}+\boldsymbol{J}\cdot\nabla\frac{\partial\mathfrak{L}}{\partial S},\end{array}\right.$ (2.24)
@@
$\varrho\,\ddot{\boldsymbol{\varphi}}=\operatorname{DIV}\big{(}\boldsymbol{P}^{\rm cons}+\boldsymbol{P}\big{)},\qquad\mathfrak{T}(\dot{S}+\operatorname{DIV}\boldsymbol{J})=\boldsymbol{P}:\nabla\dot{\boldsymbol{\varphi}}-\boldsymbol{J}\cdot\nabla\mathfrak{T},$
@@
Energy and entropy balance in the material frame.
For future comparison with the extended case later we formulate these balance laws for the general evolution equations (2.24) associated to an arbitrary Lagrangian density $\mathfrak{L}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S)$. The corresponding total energy density $\mathfrak{E}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S)$ of the continuum is defined by
$\mathfrak{E}=\frac{\partial\mathfrak{L}}{\partial\dot{\boldsymbol{\varphi}}}\cdot\dot{\boldsymbol{\varphi}}-\mathfrak{L}$
@@
Formulation with the second Piola-Kirchhoff stress tensor.
The variational formulation (2.21)–(2.23) can be equivalently expressed in terms of the second Piola-Kirchhoff stress tensor associated to $\boldsymbol{P}$ defined, as in (2.8), by
@@
Variational formulation in the spatial description.
The variational formulation in the spatial description is directly obtained by converting (2.21)–(2.23) in spatial variables. This process uses the relation (2.10) between the Lagrangian densities in material and spatial descriptions (which assumes (2.9)), as well as the relations
@@
$D_{t}f$ $=\partial_{t}f+\boldsymbol{u}\cdot\nabla f$ $D_{\delta}f=\delta f+\boldsymbol{\zeta}\cdot\nabla f$
@@
$\big{(}\boldsymbol{J}\cdot\nabla\dot{\Gamma}\big{)}\circ\boldsymbol{\varphi}^{-1}J_{\boldsymbol{\varphi}^{-1}}=\boldsymbol{j}\cdot\nabla D_{t}\gamma\quad\text{and}\quad\big{(}\boldsymbol{J}\cdot\nabla\delta\Gamma\big{)}\circ\boldsymbol{\varphi}^{-1}J_{\boldsymbol{\varphi}^{-1}}=\boldsymbol{j}\cdot\nabla D_{\delta}\gamma.$
@@
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\partial_{t}\frac{\partial\ell}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\ell}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\ell}{\partial\rho}+s\nabla\frac{\partial\ell}{\partial s}+\operatorname{div}\boldsymbol{\sigma}\\ \displaystyle-\frac{\partial\ell}{\partial s}(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\ell}{\partial s},\end{array}\right.$ (2.42)
@@
$\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}\boldsymbol{\sigma},\qquad T(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T,$
@@
Energy and entropy balance in the spatial frame.
We derive here these balance laws for an arbitrary Lagrangian density $\ell(\boldsymbol{u},\rho,s)$ for future comparison with the extended case. In this general setting the total energy density $e(\boldsymbol{u},\rho,s)$ is defined in the spatial description as
$e=\frac{\partial\ell}{\partial\boldsymbol{u}}\cdot\boldsymbol{u}-\ell.$
@@
$\frac{d}{dt}\int_{\mathcal{D}}s\,{\rm d}\mathbf{x}=\int_{\mathcal{D}}\frac{1}{T}\left(\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T\right){\rm d}\mathbf{x}-\int_{\partial\mathcal{D}}\boldsymbol{j}\cdot\boldsymbol{n}\,{\rm d}A.$
@@
$\bar{D}_{t}\varsigma=\frac{1}{T}\left(\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T\right).$
@@ 3.1 Variational formulation in the material description
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\frac{d}{dt}\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}(\dot{S}+\operatorname{DIV}\boldsymbol{J})-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\boldsymbol{J}\cdot\nabla\frac{\partial\widehat{\mathfrak{L}}}{\partial S},\end{array}\right.$ (3.4)
@@
Standard Lagrangian densities.
The natural extension of the typical Lagrangian density (2.2) to the nonequilibrium setting takes the form
@@
$\boldsymbol{P}^{\rm cons}=\varrho\frac{\partial\widehat{\mathscr{E}}}{\partial\boldsymbol{F}}\quad\text{and}\quad\mathfrak{T}=\varrho\frac{\partial\widehat{\mathscr{E}}}{\partial S}.$
@@
$\varrho\,\ddot{\boldsymbol{\varphi}}=\operatorname{DIV}\big{(}\boldsymbol{P}^{\rm cons}+\boldsymbol{P}\big{)},\qquad\mathfrak{T}(\dot{S}+\operatorname{DIV}\boldsymbol{J})=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\boldsymbol{J}\cdot\nabla\mathfrak{T}+\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}.$
Material frame indifference.
Similarly to the Lagrangian density $\mathfrak{L}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S)$ used in the reversible case and the classical irreversible case in Section 2, the nonequilibrium Lagrangian density $\widehat{\mathfrak{L}}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S,\boldsymbol{S},\boldsymbol{Q})$ is also assumed to satisfy the axiom of material indifference. As we recalled in Remark 2.1, this axiom is given by covariance of the Lagrangian density with respect to the diffeomorphisms of the ambient space $\mathbb{R}^{3}$, which act by composition on the left on $\boldsymbol{\varphi}$ and also on the ambient metric. We note that such diffeomorphisms do not act on the material fields $\varrho(t,\mathbf{X})$, $S(t,\mathbf{X})$, and on the thermodynamic fluxes $\boldsymbol{S}(t,\mathbf{X})$ and $\boldsymbol{Q}(t,\mathbf{X})$. Thus, when written in the material description, material frame indifference for $\mathfrak{L}$ and $\widehat{\mathfrak{L}}$ do not substantially differ. This is no more the case when this axiom is written in the spatial description.

### v2

@@ 2.2 Classical irreversible thermodynamics
Variational formulation in the material description. Besides the Lagrangian, in the irreversible case one needs to introduce the thermodynamic fluxes associated with the irreversible processes considered. In the case of viscosity and heat conduction the corresponding fluxes in the material description are given by the viscous Piola-Kirchhoff tensor $\boldsymbol{P}$ and the material entropy flux $\boldsymbol{J}$. In a similar way to the finite dimensional case recalled in (1.8)–(1.10), the variational formulation needs the introduction of the thermal displacement $\Gamma(t,\mathbf{X})$, see Remark 2.6, whose time rate of change is identified with the temperature in the material frame, i.e., $\frac{d}{dt}\Gamma(t,\mathbf{X})=\mathfrak{T}(t,\mathbf{X})$, as well as the additional entropy density variable $\Sigma(t,\mathbf{X})$ whose time rate of change is identified with the rate of internal entropy production of the continua $\frac{d}{dt}\Sigma(t,\mathbf{X})=\mathfrak{I}(t,\mathbf{X})\geq 0$.
@@
$\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\frac{d}{dt}\frac{\partial\mathfrak{L}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\mathfrak{L}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\mathfrak{L}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\mathfrak{L}}{\partial S}(\dot{S}+\operatorname{DIV}\boldsymbol{J})=\boldsymbol{P}:\nabla\dot{\boldsymbol{\varphi}}+\boldsymbol{J}\cdot\nabla\frac{\partial\mathfrak{L}}{\partial S},\end{array}\right$. (2.30)
@@
$\varrho\,\ddot{\boldsymbol{\varphi}}=\operatorname{DIV}\big{(}\boldsymbol{P}^{\rm cons}+\boldsymbol{P}\big{)},\qquad\mathfrak{T}(\dot{S}+\operatorname{DIV}\boldsymbol{J})=\boldsymbol{P}:\nabla\dot{\boldsymbol{\varphi}}-\boldsymbol{J}\cdot\nabla\mathfrak{T},$ (2.34)
@@
Remark 2.5 (Recovering variational formulations in the dissipationless reversible case).
There are several ways to specialize the variational formulation (2.27)–(2.29) to the reversible case.
(A) First, as expected, setting the thermodynamic fluxes to zero, i.e., $\boldsymbol{P}=0$, $\boldsymbol{J}=0$, one recovers the standard Euler-Lagrange equations for $\boldsymbol{\varphi}$, together with the entropy conservation law $\dot{S}=0$. While this variational approach may seem artificial in the absence of dissipative processes - since it still involves a variational constraint - it nevertheless demonstrates consistency with the reversible case.
(B) Second, in addition to setting $\boldsymbol{P}=0$ and $\boldsymbol{J}=0$, one may also ignore the variable $\Sigma$, and thus the associated constraints. One then simply obtains the critical point condition:
$\delta\int_{0}^{T}\!\!\int_{\mathcal{B}}\left[\mathfrak{L}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S)+S\dot{\Gamma}\right]{\rm d}^{3}\mathbf{X}\,{\rm d}t=0.$ (2.35)
In this case, entropy conservation $\dot{S}=0$ follows from the Euler-Lagrange equation associated with $\Gamma$. This setting is particularly relevant for open fluid models; see [18].
(C) Finally, in the reversible case, one may directly assume that $S$ is constant in time, as is done for $\varrho$, and omit the variables $\Gamma$ and $\Sigma$ altogether. The variational formulation (2.27)–(2.29) then reduces to the classical Hamilton principle of continuum mechanics, as given in (2.4).
Remark 2.6 (Thermal displacement).
The concept of thermal displacement $\Gamma$ plays a crucial role in the variational formulation for both classical and extended irreversible thermodynamics. However, it should be noted that $\Gamma$ is not a state variable in the heat conduction models considered in this paper, as the final equations involve only the temperature $T=\dot{\Gamma}$. Although $\Gamma$ does not appear explicitly in the final governing equations, its inclusion is essential for deriving them from a variational principle. The notion of thermal displacement was originally introduced in [65]; see also [66, 62, 35] for early discussions of this concept. This notion has since been extended to other irreversible processes beyond heat transfer, as considered in [25, 28], where the generalized variable is referred to as a thermodynamic displacement, $\Lambda_{\alpha}$, such that its time derivative $\dot{\Lambda}_{\alpha}=X_{\alpha}$ corresponds to the thermodynamic force (or affinity) $X_{\alpha}$ associated with the irreversible process $\alpha$.
Energy and entropy balance in the material frame. For future comparison with the extended case later we formulate these balance laws for the general evolution equations (2.30) associated to an arbitrary Lagrangian density $\mathfrak{L}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S)$. The corresponding total energy density $\mathfrak{E}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S)$ of the continuum is defined by
$\mathfrak{E}=\frac{\partial\mathfrak{L}}{\partial\dot{\boldsymbol{\varphi}}}\cdot\dot{\boldsymbol{\varphi}}-\mathfrak{L}$ (2.36)
@@
Formulation with the second Piola-Kirchhoff stress tensor. The variational formulation (2.27)–(2.29) can be equivalently expressed in terms of the second Piola-Kirchhoff stress tensor associated to $\boldsymbol{P}$ defined, as in (2.10), by
@@
Variational formulation in the spatial description. The variational formulation in the spatial description is directly obtained by converting (2.27)–(2.29) in spatial variables. This process uses the relation (2.12) between the Lagrangian densities in material and spatial descriptions (which assumes (2.11)), as well as the relations
@@
$D_{t}f$ $=\partial_{t}f+\boldsymbol{u}\cdot\nabla f$ $D_{\delta}f=\delta f+\boldsymbol{\zeta}\cdot\nabla f$ (2.47)
@@
$\big{(}\boldsymbol{J}\cdot\nabla\dot{\Gamma}\big{)}\circ\boldsymbol{\varphi}^{-1}J_{\boldsymbol{\varphi}^{-1}}=\boldsymbol{j}\cdot\nabla D_{t}\gamma\quad\text{and}\quad\big{(}\boldsymbol{J}\cdot\nabla\delta\Gamma\big{)}\circ\boldsymbol{\varphi}^{-1}J_{\boldsymbol{\varphi}^{-1}}=\boldsymbol{j}\cdot\nabla D_{\delta}\gamma.$ (2.52)
@@
$\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\partial_{t}\frac{\partial\ell}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\ell}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\ell}{\partial\rho}+s\nabla\frac{\partial\ell}{\partial s}+\operatorname{div}\boldsymbol{\sigma}\\ \displaystyle-\frac{\partial\ell}{\partial s}(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\ell}{\partial s},\end{array}\right$. (2.53)
@@
$\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}\boldsymbol{\sigma},\qquad T(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T,$ (2.56)
@@
Remark 2.7 (Recovering variational formulations in the dissipationless reversible case).
In parallel with Remark 2.5, we observe that the Eulerian variational formulation (2.48)–(2.51) specializes to the reversible case in three natural ways, corresponding to the Eulerian (or symmetry-reduced) analogs of the items (A)-(B)-(C) discussed in Remark 2.5.
(A) First, by setting $\boldsymbol{\sigma}=0$ and $\boldsymbol{j}=0$ in (2.48)–(2.51), one recovers the reversible fluid equations, while still involving both the phenomenological and variational constraints.
(B) Second, one may further eliminate the variable $\varsigma$, thereby removing the constraints. This leads to the critical point condition:
$\delta\int_{0}^{T}\!\!\int_{\mathcal{D}}\left[\ell(\boldsymbol{u},\rho,s)+sD_{t}\gamma\right]{\rm d}^{3}\mathbf{x}\,{\rm d}t=0,$ (2.57)
as used, e.g., in [18] for open fluid models. In this case, the entropy conservation $\bar{D}_{t}s=0$ arises as an Euler-Lagrange equation.
(C) Finally, assuming that $S$ is constant in time in the material frame directly implies $\bar{D}_{t}s=0$ in the Eulerian frame, together with the constrained variations $\delta s=-\operatorname{div}(s\boldsymbol{\zeta})$. This yields the Eulerian version of the Hamilton principle, also known as the Euler-Poincaré principle, in which the entropy density $s$ is treated analogously to the mass density $\rho$.
Energy and entropy balance in the spatial frame. We derive here these balance laws for an arbitrary Lagrangian density $\ell(\boldsymbol{u},\rho,s)$ for future comparison with the extended case. In this general setting the total energy density $e(\boldsymbol{u},\rho,s)$ is defined in the spatial description as
$e=\frac{\partial\ell}{\partial\boldsymbol{u}}\cdot\boldsymbol{u}-\ell.$ (2.58)
@@
$\frac{d}{dt}\int_{\mathcal{D}}s\,{\rm d}\mathbf{x}=\int_{\mathcal{D}}\frac{1}{T}\left(\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T\right){\rm d}\mathbf{x}-\int_{\partial\mathcal{D}}\boldsymbol{j}\cdot\boldsymbol{n}\,{\rm d}A.$ (2.62)
@@
$\bar{D}_{t}\varsigma=\frac{1}{T}\left(\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T\right).$ (2.63)
@@ 3.1 Variational formulation in the material description
$\widehat{\mathfrak{L}}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S,\boldsymbol{S},\boldsymbol{Q})$ (3.1)
@@
$\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\frac{d}{dt}\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}(\dot{S}+\operatorname{DIV}\boldsymbol{J})-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\boldsymbol{J}\cdot\nabla\frac{\partial\widehat{\mathfrak{L}}}{\partial S},\end{array}\right$. (3.5)
@@
Standard Lagrangian densities. The natural extension of the typical Lagrangian density (2.3) to the nonequilibrium setting takes the form
@@
$\boldsymbol{P}^{\rm cons}=\varrho\frac{\partial\widehat{\mathscr{E}}}{\partial\boldsymbol{F}}\quad\text{and}\quad\mathfrak{T}=\varrho\frac{\partial\widehat{\mathscr{E}}}{\partial S}.$ (3.9)
@@
$\varrho\,\ddot{\boldsymbol{\varphi}}=\operatorname{DIV}\big{(}\boldsymbol{P}^{\rm cons}+\boldsymbol{P}\big{)},\qquad\mathfrak{T}(\dot{S}+\operatorname{DIV}\boldsymbol{J})=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\boldsymbol{J}\cdot\nabla\mathfrak{T}+\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}.$ (3.10)
Material frame indifference. Similarly to the Lagrangian density $\mathfrak{L}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S)$ used in the reversible case and the classical irreversible case in Section 2, the nonequilibrium Lagrangian density $\widehat{\mathfrak{L}}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S,\boldsymbol{S},\boldsymbol{Q})$ is also assumed to satisfy the axiom of material indifference. As we recalled in Remark 2.1, this axiom is given by covariance of the Lagrangian density with respect to the diffeomorphisms of the ambient space $\mathbb{R}^{3}$, which act by composition on the left on $\boldsymbol{\varphi}$ and also on the ambient metric. We note that such diffeomorphisms do not act on the material fields $\varrho(t,\mathbf{X})$, $S(t,\mathbf{X})$, and on the thermodynamic fluxes $\boldsymbol{S}(t,\mathbf{X})$ and $\boldsymbol{Q}(t,\mathbf{X})$. Thus, when written in the material description, material frame indifference for $\mathfrak{L}$ and $\widehat{\mathfrak{L}}$ do not substantially differ. This is no more the case when this axiom is written in the spatial description.